# REA-RL: Reflection-Aware Online Reinforcement Learning for Efficient Large Reasoning Models

### v1

@@ 4.1 Online Reinforcement Learning
GRPO training.
We adopt Grouped Relative Policy Optimization (GRPO, 34). Specifically, for each question in a given dataset, GRPO samples a group of outputs in parallel from the old policy model and then optimizes the current policy model by increasing the probability of outputs with high advantage and decreasing those with low advantage, which is normalized based on the reward.
Accuracy reward.
Consistent with DeepSeek R1, we employ the rule-based accuracy reward to mitigate reward hacking. Specifically, we extract the final answer and compare it with the gold answer to verify its correctness. The reward is 1 for a correct answer and 0 otherwise.
@@
Following Kimi K1.5 [8], we incorporate a length reward to improve the LRM’s efficiency. Formally, given a group of sampled responses $\{s_{1},...,s_{G}\}$, where max_len is the length of the longest response and min_len is that of the shortest, the length reward for the $i$-th response is:
$R_{\text{Len}}(r_{i})=\left\{\begin{array}[]{rl}\lambda&\text{ if }r_{i}\text{is correct }\\ \min(0.5,\lambda)&\text{ if }r_{i}\text{ is incorrect }\end{array}\text{, where }\lambda=1-\frac{\text{len}(r_{i})-\text{min\_len}}{\text{max\_len}-\text{min\_len}}.\right.$ (1)
@@ 4.2 Reflection Model for Online Sequential Revision
In §3, we employ a 32B LLM to perform revision to detect and remove overthinking tokens through a two-step annotation process. To achieve better scalability, we train a smaller reflection model, effectively providing sequential revision online, thus offering an additional dimension of scaling beyond traditional parallel sampling.
@@
To construct SFT data, we generate four responses per question on the training dataset using R1-7B, retaining only correct responses. Each chunk within these responses is labeled based on whether it contains the correct answer following §3.1. We then construct the SFT data using the question and all response chunks as input, training the reflection model to predict the category of each chunk. For efficiency, we adopt a concise output format, directly generating the ID and its category without additional deliberation, as detailed in Appendix A.
@@
As depicted in Figure 2, following the parallel sampling of multiple responses in online RL, we utilize the reflection model for sequential revision. Specifically, for the sampled response $S=\{s_{1},...,s_{G}\}$, we first use the reflection model $\text{M}_{\text{Reflect}}$ to remove overthinking tokens (red segment), copies the preceding segments from the original responses $S$ (yellow segment), and prompt the policy model $\text{M}_{\text{Policy}}$ to generate a final answer (blue segment), which forms the revision $S^{r}=\{s_{1}^{r},...,s_{G}^{r}\}$. Cases are shown in Appendix B.1. Ultimately, both the original responses $S$ and the revised responses $S^{r}$ are used as training data for the policy model, enabling an additional dimension of scaling and guiding the optimization direction of RL. Formally:
[…]
Equivalence to a partial reward.
Given the substantial similarity between the pre- and post-revision responses (highlighted in blue), we can decompose the reward into two components. Formally, let $a_{i}$ and $a_{i}^{r}$ represent the advantages of the original response $s_{i}$ and revised response $s_{i}^{r}$, respectively. The first component is the average reward across all tokens in both responses, formulated as $(a_{i}^{r}+a_{i})/2$. The second component is a partial reward. Specifically, the overthinking tokens (red) receive a reward of $-(a_{i}^{r}-a_{i})/2$, while the revised tokens (blue) receive $(a_{i}^{r}-a_{i})/2$. When revision is successful, the length reward ensures $a_{i}^{r}>a_{i}$, effectively penalizing overthinking.
However, if both the original and revised responses are incorrect, it suggests a need for more extensive reasoning, making a preference for shorter outputs detrimental. Furthermore, if revision transforms a correct answer into an incorrect one, the opposite effect occurs, potentially encouraging overthinking. To mitigate these issues, we discard the revised responses in such cases and retain the original ones. Finally, following Yu et al. [45], when all generated trajectories yield incorrect answers, we skip these instances to avoid unintended optimization.
@@ 4.3 Reflection-Aware Reward Refinement
To prevent the models from favoring short yet non-reflective responses due to length reward, we introduce a reflection reward based on the presence of reflective tokens, i.e., “wait”, “alternatively”, “check”, and “but”. We calculate the density of these tokens, and count closely clustered occurrences as a single instance to prevent consecutive reflective tokens within a single reflection. Formally, for the current response $s_{i}$, let $N_{\text{Token}}$ be the total number of tokens in the response and $N_{\text{Reflect}}$ be the number of reflective tokens. The reflection reward $R_{\text{Reflect}}$ is:
$R_{\text{Reflect}}(s_{i})=\min(0,\frac{D_{i}}{D_{0.2}}-1)\text{, where }D_{i}=\frac{N_{\text{Reflect}}}{N_{\text{Token}}}.$ (3)
$D_{0.2}$ represents the 0.2 quantile of the reflection density in the training data, and $D_{i}$ is the density of $r_{i}$. This reward penalizes responses only when their density falls within the lowest 20% of observed densities, thereby preventing the reward from inadvertently promoting overthinking.
@@
The length reward in Kimi K1.5 demonstrates a bias towards shorter responses even if the answers are incorrect. However, for challenging queries, more extensive reasoning is often required, conflicting with the current reward mechanism. Therefore, we follow Zhang and Zuo [46] to refine the length reward by grouping responses by query and setting the length reward to zero for all incorrect responses, which provides the definition of $R_{\text{RLen}}$.
@@ 5.1 Experimental Setup
Common training configuration.
Unless explicitly stated otherwise, we maintain the following experimental settings. We use the DeepScaleR 40k dataset [25] as the training data and DeepSeek-R1-Distill-Qwen-7B (R1-7B) as the base model. We only retain questions for which the model can provide at least one correct answer within 4 samples as the training data, resulting in 30k questions. We train for 1800 steps across all methods and baselines, with each batch containing 16 different questions and 4 different paths generated for each question, with a maximum generation length of 8k. We use a learning rate of $2\times 10^{-5}$, integrating low-rank adaptation [14] with all attention and mlp parameters, setting the LoRA rank $r$ to 16 and alpha to 32. During the data generation process, we follow R1 to use a temperature of 0.6 and a top_p of 0.95. Finally, the reflection model is fine-tuned from Qwen2.5-7B-Instruct for one epoch. During sequential revision, we set the temperature to 0 for the reflection model.
@@
$\text{M}_{\text{Reflect}}$” represents online revision in §4.2. “$\text{R}_{\text{Len}}$” represents adding the Kimi K1.5 length reward, “$\text{R}_{\text{RLen}}$” represents adding our refined length reward, and “$\text{R}_{\text{Reflect}}$” represents adding the reflection reward in §4.3. Accuracy reward is used in all GRPO settings. The primary baselines are 1) GRPO, which uses accuracy reward only. 2) GRPO $\text{R}_{\text{Len}}$, which uses accuracy reward along with the Kimi K1.5 length reward.
@@
We also compare with commonly used offline baselines. We use the revised responses generated from the 32B model as the training data, following §3, with other generation settings consistent with the online approach. Only data that the answer is correct after revision is used for training. We then conduct training using 1) SFT [47]: We simply fine-tune the model using the aforementioned revised dataset. 2) RPO [30]: We treat the revised response as positive examples and the original paths as negative examples.
@@ 5.2 Main Result
Method GSM8K Math500 Gaokao23 Amc23 Olympiad Aime24 Average

### v2

@@ 4.1 Online Reinforcement Learning
Online RL.
We adopt Grouped Relative Policy Optimization (GRPO, Shao et al., 2024). Specifically, for each question in a given dataset, GRPO samples a group of paths $S=\{s_{1},...,s_{G}\}$ in parallel from the policy model. Then, for each path $s_{i}$, we calculate its reward $r_{i}$ as the sum of several reward functions introduced later. These rewards are normalized within the group $S$ to get the advantage of each path, i.e., $a_{i}=\frac{r_{i}-\operatorname{mean}(\mathbf{r})}{\operatorname{std}(\mathbf{r})}$. Finally, the policy model is optimized by increasing the probability of paths with high advantage and decreasing the probability of those with low advantage.
Following Shao et al. (2024), we apply the rule-based accuracy reward $\text{R}_{\text{Acc}}$ to mitigate reward hacking. Specifically, we extract the final answer and compare it with the gold answer to verify its correctness. The reward is 1 for a correct answer and 0 otherwise.
@@
Following Kimi K1.5 (Du et al., 2025), we incorporate a length reward to improve efficiency. Formally, given a group of sampled responses $\{s_{1},...,s_{G}\}$, where max_len is the length of the longest response and min_len is that of the shortest, the length reward for the $i$-th response is:
[…]
Figure 2: Workflow of REA-RL. We first parallel sample $G$ paths as in GRPO. Then, the reflection model identifies and truncates overthinking tokens (red ones with $\times$), retaining the preceding yellow segments. After that, the policy model finishes the truncated paths, generating revised tokens (blue ones with $\checkmark$), and yielding $G$ revised paths. Finally, both the $G$ original and $G$ revised paths are used for training. Cases are in Appendix B.1. In addition to the naive accuracy reward $\text{R}_{\text{Acc}}$ and length reward $\text{R}_{\text{Len}}$, we refine the length reward ($\text{R}_{\text{RLen}}$) and introduce a new reflection reward ($\text{R}_{\text{Reflect}}$).
@@ 4.2 Reflection Model for Online Sequential Revision
Design principle.
We use a reflection model to produce shorter online revisions of the trained model’s reasoning path, using the same revision definition as §3.1. The task is simple, but enforcing per-sentence judgments over long traces is challenging. As evaluated in §3.2, a 32B LLM performs well with a two-step annotation, but is too slow. Qwen2.5-7B-Instruct (Qwen-7B) is faster, yet in 17% of cases its outputs violate the required judgment count. Given the task’s simplicity, we distill the 32B two-step revision ability into Qwen-7B with SFT and run it in one step to ensure efficiency.
@@
To construct SFT data, we generate four responses per question on the training dataset using R1-7B, retaining only correct responses. Each chunk within these responses is labeled based on whether it contains the correct answer following §3.1. We then construct the SFT data using the question and all response chunks as input, training the reflection model to predict the category of each chunk in a single step. For efficiency, we adopt a concise output format, directly generating the ID and its category without additional deliberation, as detailed in Appendix A.
@@
As depicted in Figure 2, following the parallel sampling of multiple responses in online RL, we utilize the reflection model for sequential revision. Specifically, for the sampled responses $S=\{s_{1},...,s_{G}\}$, we first use the reflection model $\text{M}_{\text{Reflect}}$ to remove overthinking tokens (red segment), copy the preceding segments from the original responses $S$ (yellow segment), and prompt the policy model $\text{M}_{\text{Policy}}$ to generate a final answer (blue segment), which forms the revision $S^{r}=\{s_{1}^{r},...,s_{G}^{r}\}$. Cases are shown in Appendix B.1. Ultimately, both the original responses $S$ and the revised responses $S^{r}$ are used as training data for the policy model, enabling an additional dimension of scaling and guiding the optimization direction of RL. Formally:
$\{s_{1},...,s_{G},s_{1}^{r},...,s_{G}^{r}\}\xrightarrow{\textbf{Online RL}}\mathbf{\text{M}_{\text{Policy}}}\text{, where }s_{i}^{r}=\text{M}_{\text{Policy}}(\text{M}_{\text{Reflect}}(s_{i})).$ (2)
Equivalence to a partial advantage.
We find that our method is equivalent to a partial penalty only for the overthinking portion. Given the substantial similarity between the pre- and post-revision responses, we can decompose the advantage into two components. Let $a_{i}$ and $a_{i}^{r}$ represent the advantages of the original response $s_{i}$ and revised response $s_{i}^{r}$, respectively. The first component is the average advantage for all tokens in both responses, formulated as $(a_{i}^{r}+a_{i})/2$. The second component is a partial advantage for the different parts. The overthinking tokens (red) receive a penalty of $(a_{i}^{r}-a_{i})/2$, while the revised tokens (blue) receive $(a_{i}^{r}-a_{i})/2$. When revision is successful, the length reward ensures $a_{i}^{r}>a_{i}$, effectively penalizing overthinking.
However, if revision transforms a correct answer into an incorrect one, the opposite effect occurs, potentially encouraging overthinking. Furthermore, if both the original and revised responses are incorrect, it suggests a need for more extensive reasoning, making a preference for shorter outputs detrimental. Therefore, we discard the revised responses in such cases and retain the original ones.
@@ 4.3 Reflection-Aware Reward Refinement
To prevent the models from favoring short yet non-reflective responses due to the length reward, we introduce a reflection reward based on the presence of reflective tokens, i.e., “wait”, “alternatively”, “check”, and “but”. We calculate the density of these tokens, and count closely clustered occurrences as a single instance to prevent consecutive reflective tokens within a single reflection. Formally, for the current response $s_{i}$, let $N_{\text{Token}}$ be the total number of tokens in the response and $N_{\text{Reflect}}$ be the number of reflective tokens. The reflection reward $\text{R}_{\text{Reflect}}$ is:
$\text{R}_{\text{Reflect}}(s_{i})=\min(0,\frac{D_{i}}{D_{0.2}}-1)\text{, where }D_{i}=\frac{N_{\text{Reflect}}}{N_{\text{Token}}}.$ (3)
Here, $D_{0.2}$ represents the 0.2 quantile of the reflection density in the training data, and $D_{i}$ is the density of $s_{i}$. This reward penalizes responses only when their density falls within the lowest 20% of observed densities, thereby preventing the reward from inadvertently promoting overthinking. Empirically, using any quantile $\leq$ 0.2 yields similarly strong performance; a smaller quantile places greater emphasis on efficiency, whereas a larger one trades efficiency for accuracy.
@@
The length reward in Kimi K1.5 demonstrates a bias towards shorter responses even when incorrect. However, for challenging queries, more extensive reasoning is often required, conflicting with the current reward mechanism. Thus, we follow Zhang and Zuo (2025) to set the length reward to zero if the response is incorrect, i.e.:
[…]
@@ 5.1 Experimental Setup
We maintain most settings across experiments, with training configs in Appendix B.2. Baselines are:
@@
$\text{M}_{\text{Reflect}}$” represents online revision in §4.2. $\text{R}_{\text{Len}}$” represents adding the Kimi K1.5 length reward, $\text{R}_{\text{RLen}}$” represents adding our refined length reward, and $\text{R}_{\text{Reflect}}$” represents adding the reflection reward in §4.3. Accuracy reward is used in all GRPO settings. The primary baselines are 1) GRPO, which uses accuracy reward only. 2) GRPO $\text{R}_{\text{Len}}$, which uses accuracy reward along with the Kimi K1.5 length reward.
@@
We also compare with commonly used offline baselines. We use the revised responses generated from the 32B model as the training data, following §3, with other generation settings consistent with the online approach. Only data that the answer is correct after revision is used for training. We then conduct training using 1) SFT (Zhang et al., 2023): We simply fine-tune the model using the aforementioned revised dataset. 2) RPO (Pang et al., 2024): We treat the revised responses as positive examples and the original responses as negative examples.
For further comparison, we evaluate against related methods initialized from R1-7B, including DAST (Shen et al., 2025), Light-R1 (Wen et al., 2025), ShorterBetter (Yi et al., 2025), and Arora and Zanette (2025). We download their publicly available checkpoints and run the same evaluation. For the prompt-based method NoThink (Ma et al., 2025), we use its released prompt configuration.
Evaluation dataset.
We follow Yang et al. (2024b) in evaluating our approach on five math test sets, ordered by difficulty: GSM8K (Cobbe et al., 2021), 8.5K grade school math problems; MATH500 (Hendrycks et al., 2021b), 500 challenging high school competition problems; Gaokao23 (Liao et al., 2024), English-translated math questions from the 2023 Chinese Gaokao exam; Amc23111https://huggingface.co/datasets/AI-MO/aimo-validation-amc, 2023 American Mathematics Competitions; and Aime24222https://huggingface.co/datasets/AI-MO/aimo-validation-aime, 2024 American Invitational Mathematics Examination. We employ the math validator provided by rStar (Guan et al., 2025). Considering the limited size of Amc23 and Aime24, we sample 8 paths for each question to mitigate randomness. We report additional experimental results on more models and evaluation domains in Appendix B.3.
Reflection model training.
We employ R1-7B to generate reasoning processes on DeepScaleR, sampling four paths for each question. In 74.03% of the questions, the model generates at least one correct answer in 8k tokens, and in 46.54% of the questions, all four answers are correct. Subsequently, we label the paths with correct answers as in §3.1 to form the training data. We then fine-tune Qwen2.5-7B-Instruct for one epoch. During online sequential revision, we set the temperature to 0.
Method GSM8K Math500 Gaokao23 Amc23 Aime24 Average
[…]
Budget: 8k Original 91.58 1001312 88.40 1003389 77.40 1003270 77.50 1004746 40.83 1006815 75.14 1003906
SFT 89.92 64.63 89.20 69.25 77.92 72.91 82.81 72.67 41.25 89.98 76.22 73.89
RPO 90.14 51.22 89.80 53.35 83.12 55.20 80.62 59.54 41.67 74.78 77.07 58.82
GRPO 92.80 117.99 89.00 100.74 79.22 101.31 79.69 96.23 39.17 98.88 75.98 103.03
GRPO $\text{R}_{\text{Len}}$ (Du et al., 2025) 85.97 32.01 87.60 59.04 70.91 58.84 85.62 71.43 45.42 87.42 75.10 61.75
NoThink (Ma et al., 2025) 85.06 20.50 83.20 27.94 65.71 25.81 65.00 28.82 20.42 44.37 63.88 29.49
ShorterBetter (Yi et al., 2025) 85.37 15.17 83.40 29.30 66.75 25.96 76.56 43.43 48.33 68.67 72.08 36.51
Light-R1 (Wen et al., 2025) 86.28 39.25 87.00 77.40 68.05 80.89 75.31 97.01 42.92 96.92 71.91 78.29
DAST (Shen et al., 2025) 87.79 34.60 89.20 77.49 81.30 78.62 84.06 87.40 42.50 96.71 76.97 74.96
Arora and Zanette (2025) 90.45 40.32 91.20 73.21 76.10 70.55 82.50 84.68 43.33 94.34 76.72 72.62
GRPO $\text{R}_{\text{Len}}$ $\text{M}_{\text{Reflect}}$ 89.23 37.80 91.20 48.66 78.44 52.23 86.56 59.65 41.67 84.15 77.42 56.50
GRPO $\text{R}_{\text{RLen+Reflect}}$ 92.72 67.07 91.40 69.52 80.52 72.54 85.62 75.64 47.92 91.37 79.64 75.23
GRPO $\text{R}_{\text{RLen+Reflect}}$ $\text{M}_{\text{Reflect}}$ 89.99 53.12 89.60 61.79 81.30 61.19 85.62 71.39 42.92 88.45 77.89 67.19
Budget: 16k Original 91.66 1001344 92.00 1003893 81.82 1003785 88.12 1005840 48.33 10010460 80.39 1005064
SFT 89.99 68.75 90.60 69.25 79.48 74.37 88.44 72.11 48.75 89.68 79.45 74.83
RPO 90.14 50.00 89.80 46.90 83.12 47.77 82.19 50.75 42.92 51.41 77.63 49.37
GRPO 92.87 116.52 93.40 99.20 82.34 100.26 87.19 97.31 50.42 97.18 81.24 102.09
GRPO $\text{R}_{\text{Len}}$ (Du et al., 2025) 85.97 31.25 88.40 56.43 72.21 55.88 87.81 65.65 50.00 76.96 76.88 57.23
NoThink (Ma et al., 2025) 85.06 20.01 84.20 27.23 66.23 24.70 66.25 26.10 23.33 38.38 65.01 27.28
ShorterBetter (Yi et al., 2025) 85.37 14.81 83.40 27.02 66.75 22.85 76.88 37.52 50.00 51.59 72.48 30.76
Light-R1 (Wen et al., 2025) 86.28 38.32 91.60 77.40 74.55 81.88 88.75 96.73 54.17 88.15 79.07 76.50
DAST (Shen et al., 2025) 87.79 34.67 91.40 79.84 83.12 80.79 88.44 87.12 51.67 98.41 80.48 76.17
Arora and Zanette (2025) 90.45 39.58 93.20 70.15 77.14 70.20 86.88 83.39 48.33 90.25 79.20 70.71
GRPO $\text{R}_{\text{Len}}$ $\text{M}_{\text{Reflect}}$ 89.23 36.90 92.40 45.11 79.74 47.50 88.12 56.04 47.92 75.82 79.48 52.27
GRPO $\text{R}_{\text{RLen+Reflect}}$ 92.72 65.48 92.80 66.71 81.82 68.27 88.75 72.17 54.58 86.21 82.13 71.77
GRPO $\text{R}_{\text{RLen+Reflect}}$ $\text{M}_{\text{Reflect}}$ 89.99 52.08 91.00 60.78 82.08 55.85 89.38 67.76 51.25 81.09 80.74 63.51
Table 2: Main results of our proposed methods. Most abbreviations align with Table 1. Baseline definitions are in §5.1. “GRPO $\text{R}_{\text{Len}}$ $\text{M}_{\text{Reflect}}$” represents the addition of the reflection model, “GRPO $\text{R}_{\text{RLen+Reflect}}$” represents the addition of the reward optimization, and “GRPO $\text{R}_{\text{RLen+Reflect}}$ $\text{M}_{\text{Reflect}}$” represents the combination of both optimizations. “Budget” is the max tokens allowed per question.
@@ 5.2 Main Result
Improved performance compared to offline method.
The results are in Table 2. The final three rows showcase the performance of the reflection model, reward refinement, and their combined approach. While RPO shares similar ideas with the sequential revision, the primary distinction lies in the offline dataset. RPO achieves comparable performance at an 8k budget but significantly degrades on two more challenging datasets at a 16k budget. In contrast, our method demonstrates better performance, highlighting the advantage of online training over offline training.
GRPO, using only an accuracy reward, shows no significant performance improvement, suggesting that gains are not solely due to more training. Adding a length reward greatly reduces inference costs but severely hurts performance. Relative to all other baselines including GRPO $\text{R}_{\text{Len}}$, we consistently obtain higher accuracy at comparable truncation ratios. While some baselines achieve notably more aggressive truncation, they incur large accuracy drops that undermine their practicality.
Reflection model and reflection reward target distinct dimensions.
The reflection model is more effective in reducing response length, whereas reflection reward contributes more to accuracy enhancement. By integrating both strategies, a balanced outcome can be achieved, yielding a 36% efficiency improvement without compromising performance. However, the improvement under an 8k budget remains limited. This can be attributed to the advantage of shorter generations under a constrained budget, and the reflection model’s inherent ability to foster reflection, as analyzed in §5.4. The effectiveness of $\text{R}_{\text{RLen}}$ and $\text{R}_{\text{Reflect}}$ and the choice of quantile is verified in Appendix B.4.
5.3 Reflection Model Analysis
Experimental setup.
We follow §3.2 to verify our reflection model $\text{M}_{\text{Reflect}}$. We use $\text{M}_{\text{Reflect}}$ to truncate paths generated by R1-7B, then have R1-7B produce the final answer. The truncation is validated by the correctness of this answer. We compare against the untrained 7B and 32B LLMs and a fixed truncation strategy. For $\text{M}_{\text{Reflect}}$, we define three revision strengths: Normal truncates at the first identified correct answer; Weak truncates at the second such position; and Strong truncates before the first position where the truncation probability exceeds 0.25. If no such position exists, the Normal position is used. Fixed Trunc truncates at the closest sentence-ending position matching the truncation ratio achieved by our method, and then prompts R1-7B to generate the final answer.
Comparable performance to 32B at a significantly lower cost.
Results are shown in Table 3. Our reflection model does not incorporate the gold answer as input. It significantly outperforms its 7B base model (7B Revise). In comparison to the 32B model revised without gold input, our Strong strategy achieves a comparable compression ratio while exhibiting minimal performance degradation, demonstrating that our model achieves comparable results at a substantially reduced cost. Furthermore, our method also outperforms fixed truncation, especially when the truncation ratio is large. Finally, we observe that under certain evaluations, the accuracy actually increases after revision. This indicates that the model may have already arrived at the correct answer, but fails to explicitly state this conclusion under the 16k token budget.
REA-RL outperforms inference with reflection model.
To ensure revision accuracy, we employ the Normal strategy during training. However, REA-RL with the reflection model achieves a higher compression ratio than the Strong strategy while maintaining comparable performance, despite the increased inference cost associated with the generate-then-revise approach of the Strong strategy. This improvement can be attributed to the length reward and the online training, which iteratively train the model and provide further guidance for shortening already concise responses.
Online revision as an efficient scaling strategy.
The reflection model provides revised responses, which doubles the dataset size. To evaluate whether other methods could yield similar benefits, we extend the GRPO baseline generation to 8 paths, i.e., “Gen8”, to verify whether scaling parallel sampling is more effective. However, Gen8 indicates no consistent improvement over parallel sampling of 4 paths, while our combination with sequential revision, under the same reward $\text{R}_{\text{Len}}$, demonstrates superior performance and truncation ratios, thereby establishing the enhanced efficacy of our scaling method compared to mere parallel sampling.
Regarding training time, we optimize the implementation by deferring the update of the vllm inference model, which enables parallel execution of training and data generation, resulting in approximately a twofold speedup. On 3 NVIDIA A800 80G GPUs, GRPO with sampling 4 paths requires 80 hours, while sampling 8 paths requires 110 hours. Our method, incorporating the reflection model, requires 120 hours. Given the significant performance improvement, this additional computational cost is deemed acceptable. Furthermore, our approach actually reduces token usage and can lower cost in large-scale experiments, detailed in Appendix B.5. Finally, if training efficiency is a primary concern, utilizing the reflection reward alone also requires only 80 hours and achieves improved performance.
Method GSM8K Math500 Gaokao23 Amc23 Aime24 Average